OpenStudy (anonymous):

plz helppp: what value of x in the interval 0 degrees less than and equal to X is less then and equal to 180 degrees satisfies the equation radical3 tanx+1=0?

OpenStudy (anonymous):

\[0 \le x \le 180\]

OpenStudy (anonymous):

\[\sqrt{3}\tan x+1=0\]

OpenStudy (mathmale):

Hello, lyubas! Please solve this latest equation for tan x. Then we'll solve for x itself.

OpenStudy (anonymous):

do u subtract 1 on both sides?

OpenStudy (mathmale):

Yes, subtract 1 from both sides. Don't be afraid to make mistakes...jump in and do whatever you think is appropriate at the moment.

OpenStudy (anonymous):

I got\[\sqrt{3} \tan x=-1\]

OpenStudy (anonymous):

can I divde by radical 3 on both sides?

OpenStudy (mathmale):

Nice. Now solve for tan x by dividing both sides by Sqrt(3).

OpenStudy (anonymous):

i got -.5773502692

OpenStudy (mathmale):

I'd rather that you write \[\tan x=\frac{ -1 }{ \sqrt{3} }\]

OpenStudy (anonymous):

oh ok

OpenStudy (mathmale):

In which quadrants is it possible for the tangent to be negative?

OpenStudy (anonymous):

2,4

OpenStudy (mathmale):

Right. So your angle (x) could be in either Q2 or Q4. Let's assume for the moment that x is in Q2. What is the angle? Draw a picture. Note that the angle would have to be between 90 and 180 degrees to be in Q4.

OpenStudy (anonymous):